# SAT Encoding of Partial Ordering Models for Graph Coloring Problems

### v1

@@ 4.4 Experimental results for the bandwidth coloring problem
set #inst. ASS-S-B POP-S-B POPH-S-B ASS-I-B POP-I-B POPH-I-B DFMM
@@ 5 Conclusion
In this paper, we have suggested new partial-ordering based ILP and SAT formulations for graph coloring and bandwidth coloring problems.
Our computational study on the graph coloring problem suggests that all three SAT encodings perform similar, with (POP-S) and (POPH-S) solving slightly more DIMACS instances. This holds true for every timelimit up to 1 hour. Moreover, the SAT encodings solve more instances than the ILP formulations. Compared to state-of-the-art approaches, the tested SAT based approaches solved more instances than the approach based on the set cover ILP formulations and have shown to be particularly advantageous for sparse graphs. Moreover, the tested SAT based encodings also solved more instances than reported in [7]. Specifically, (POP-S) and (POPH-S) have proven to be effective on the wap0-instances
@@ Appendix A Detailed results of the nine models for the 143 DIMACS instances for the GCP
POP-S POPH-S ASS-S POP-I POPH-I ASS-I EC CDCL CLICOL
@@
1-FullIns_3 30 100 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 0.0 0 0
1-FullIns_4 93 593 5 5 0.1 5 5 0.1 5 5 0.1 5 5 0.1 5 5 0.2 5 5 0.1 3600.1 0 0
1-FullIns_5 282 3247 6 6 0.5 6 6 0.6 6 6 0.5 6 6 361.2 6 6 242.9 6 6 29.4 3600.5 0 0
1-Insertions_4 67 232 5 5 1.4 5 5 1.2 5 5 1.0 5 5 71.7 5 5 480.4 5 5 106.9 3600.1 3600.0 2
@@
2-FullIns_3 52 201 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 0.0 0 0
2-FullIns_4 212 1621 6 6 0.0 6 6 0.0 6 6 0.1 6 6 1.8 6 6 0.4 6 6 0.2 3600.1 0 0
2-FullIns_5 852 12201 7 7 0.7 7 7 0.6 7 7 0.8 6 7 3600.0 7 7 1899.5 7 7 983.6 3602.5 20 7
2-Insertions_3 37 72 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.1 4 4 0.1 4 4 0.1 220.2 0 0
@@
3-FullIns_3 80 346 6 6 0.1 6 6 0.1 6 6 0.1 6 6 0.0 6 6 0.0 6 6 0.0 0.0 0 0
3-FullIns_4 405 3524 7 7 0.1 7 7 0.1 7 7 0.1 7 7 7.6 7 7 1.4 7 7 0.4 3600.2 0 1
3-FullIns_5 2030 33751 8 8 4.1 8 8 5.9 8 8 5.8 7 8 3600.0 7 8 3600.0 7 8 3600.0 3602.9 0 93
3-Insertions_3 56 110 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.8 4 4 1.1 4 4 1.4 3600.1 1 0
@@
4-FullIns_3 114 541 7 7 0.1 7 7 0.1 7 7 0.0 7 7 0.0 7 7 0.0 7 7 0.0 0.0 0 0
4-FullIns_4 690 6650 8 8 0.4 8 8 0.5 8 8 0.6 8 8 60.1 8 8 0.6 8 8 0.5 3600.7 0 2
4-FullIns_5 4146 77305 9 9 63.8 9 9 93.3 9 9 75.6 8 9 3600.0 8 9 3600.0 8 9 3600.0 3629.8 730 2137
4-Insertions_3 79 156 4 4 0.1 4 4 0.1 4 4 0.2 4 4 17.5 4 4 15.7 4 4 48.3 3600.2 417 0
@@
anna 138 493 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 0.0 0 0
ash331GPIA 662 4181 4 4 0.1 4 4 0.1 4 4 0.1 4 4 6.3 4 4 16.0 4 4 28.6 3611.7 0 0
ash608GPIA 1216 7844 4 4 0.1 4 4 0.1 4 4 0.0 4 4 10.7 4 4 34.1 4 4 74.3 3600.0 4 1
ash958GPIA 1916 12506 4 4 0.1 4 4 0.1 4 4 0.1 4 4 57.5 4 4 217.9 4 4 872.6 3600.0 29 2
david 87 406 11 11 0.1 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 0.0 0 0
@@
DSJR500.1 500 3555 12 12 0.0 12 12 0.0 12 12 0.0 12 12 1.2 12 12 0.3 12 12 0.2 255.1 0 0
DSJR500.1c 500 121275 81 89 3600.0 80 89 3600.0 80 89 3600.0 74 0 3600.1 81 0 3600.0 75 89 3600.3 695.1 3600.0 123
DSJR500.5 500 58862 122 131 3600.0 122 131 3600.0 122 131 3600.0 115 0 3600.1 122 122 334.5 122 122 1373.1 2602.2 3600.0 3600.0
@@
fpsol2.i.1 269 11654 65 65 0.1 65 65 0.1 65 65 0.1 65 65 1.2 65 65 0.3 65 65 0.3 0.4 0 0
fpsol2.i.2 363 8691 30 30 0.1 30 30 0.2 30 30 0.0 28 30 3600.0 30 30 0.2 30 30 0.4 0.4 0 0
fpsol2.i.3 363 8688 30 30 0.0 30 30 0.0 30 30 0.0 29 30 3600.0 30 30 0.2 30 30 0.4 0.4 0 0
games120 120 638 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 0.0 0 0
homer 556 1629 13 13 0.0 13 13 0.0 13 13 0.0 13 13 61.2 13 13 0.1 13 13 0.1 0.1 0 0
huck 74 301 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.2 11 11 0.0 11 11 0.0 0.0 0 0
inithx.i.1 519 18707 54 54 0.1 54 54 0.0 54 54 0.1 54 54 0.8 54 54 0.3 54 54 0.3 1.3 0 1
inithx.i.2 558 13979 31 31 0.1 31 31 0.0 31 31 0.1 29 31 3600.0 31 31 0.2 31 31 0.4 0.2 0 0
inithx.i.3 559 13969 31 31 0.1 31 31 0.2 31 31 0.1 29 31 3600.0 31 31 0.2 31 31 0.3 0.3 0 0
jean 77 254 10 10 0.0 10 10 0.1 10 10 0.0 10 10 0.0 10 10 0.0 10 10 0.0 0.0 0 0
latin_square_10 900 307350 90 132 3600.0 90 132 3600.0 90 132 3600.0 - - - - - - - - - 3638.4 3600.0 3600.0
le450_15a 450 8168 15 15 0.8 15 15 0.3 15 15 0.4 15 16 3600.1 15 15 328.1 15 15 307.6 3633.7 4 1
le450_15b 450 8169 15 15 0.7 15 15 0.2 15 15 0.2 15 15 1039.7 15 15 567.0 15 15 293.2 3600.0 1 0
le450_15c 450 16680 15 23 3600.0 15 23 3600.0 15 23 3600.0 15 23 3600.0 15 0 3600.0 15 0 3600.7 3600.0 3600.0 212
le450_15d 450 16750 15 24 3600.0 15 24 3600.0 15 24 3600.0 15 0 3600.0 15 0 3600.0 15 0 3600.4 3600.0 3600.0 3600.0
le450_25a 450 8260 25 25 0.1 25 25 0.1 25 25 0.1 25 25 26.8 25 25 1.3 25 25 0.6 1.8 0 0
le450_25b 450 8263 25 25 0.0 25 25 0.0 25 25 0.0 25 25 10.9 25 25 0.8 25 25 0.6 1.9 0 0
Table 3: Results of the nine models for the 143 DIMACS instances for the GCP
POP-S POPH-S ASS-S POP-I POPH-I ASS-I EC CDCL CLICOL
@@
le450_5a 450 5714 5 5 0.1 5 5 0.0 5 5 0.0 5 5 108.5 5 5 23.2 5 5 62.9 3600.0 41 0
le450_5b 450 5734 5 5 0.0 5 5 0.0 5 5 0.0 5 5 328.8 5 5 41.2 5 5 47.6 3600.0 9 1
le450_5c 450 9803 5 5 0.0 5 5 0.1 5 5 0.0 5 5 92.4 5 5 47.5 5 5 19.3 3600.0 2 0
le450_5d 450 9757 5 5 0.1 5 5 0.1 5 5 0.0 5 5 19.8 5 5 49.8 5 5 46.8 3584.7 2 0
miles1000 128 3216 42 42 0.1 42 42 0.1 42 42 0.1 42 42 0.5 42 42 0.1 42 42 0.1 0.1 0 0
miles1500 128 5198 73 73 0.1 73 73 0.2 73 73 0.2 73 73 0.5 73 73 0.3 73 73 0.3 0.1 0 0
miles250 125 387 8 8 0.1 8 8 0.0 8 8 0.0 8 8 0.0 8 8 0.0 8 8 0.0 0.0 0 0
miles500 128 1170 20 20 0.1 20 20 0.0 20 20 0.1 20 20 0.0 20 20 0.0 20 20 0.0 0.0 0 0
miles750 128 2113 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.1 31 31 0.1 31 31 0.1 0.0 0 0
mug100_1 100 166 4 4 0.1 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.1 4 4 0.2 0.6 0 0
mug100_25 100 166 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.4 4 4 0.2 0.5 0 0
mug88_1 88 146 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.1 4 4 0.1 4 4 0.2 0.3 0 0
mug88_25 88 146 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.1 4 4 0.2 0.3 0 0
mulsol.i.1 138 3925 49 49 0.0 49 49 0.1 49 49 0.1 49 49 0.4 49 49 0.1 49 49 0.1 0.1 0 0
mulsol.i.2 173 3885 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.3 31 31 0.1 31 31 0.1 0.0 0 1
mulsol.i.3 174 3916 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.2 31 31 0.1 31 31 0.1 0.0 0 0
mulsol.i.4 175 3946 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.4 31 31 0.1 31 31 0.1 0.0 0 0
mulsol.i.5 176 3973 31 31 0.0 31 31 0.0 31 31 0.0 30 31 3600.0 31 31 0.1 31 31 0.1 0.0 0 0
myciel3 11 20 4 4 0.0 4 4 0.1 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 0.0 0 0
myciel4 23 71 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.1 5 5 0.1 5 5 0.1 4.8 0 0
myciel5 47 236 6 6 0.2 6 6 0.2 6 6 0.2 6 6 37.5 6 6 42.0 6 6 44.3 3600.0 0 1
myciel6 95 755 7 7 99.4 7 7 74.3 7 7 63.6 6 7 3600.0 6 7 3600.0 5 7 3600.0 3600.1 0 1422
myciel7 191 2360 6 8 3600.0 6 8 3600.0 6 8 3600.0 5 8 3600.0 5 8 3600.0 5 8 3600.0 3600.7 0 3600.0
@@
qg.order30 900 26100 30 30 0.8 30 30 0.4 30 30 0.5 30 35 3600.2 30 30 77.6 30 30 186.5 3600.0 0 11
qg.order40 1600 62400 40 40 7.5 40 40 1.9 40 40 2.4 40 0 3600.0 40 40 1798.7 40 43 3600.2 3600.0 8 228
qg.order60 3600 212400 60 60 1813.9 60 60 835.9 60 60 28.8 60 0 3600.1 60 62 3601.5 60 0 3600.0 3600.0 347 3600.0
@@
queen5_5 25 160 5 5 0.0 5 5 0.0 5 5 0.1 5 5 0.0 5 5 0.0 5 5 0.0 0.0 0 0
queen6_6 36 290 7 7 0.1 7 7 0.1 7 7 0.0 7 7 0.4 7 7 0.2 7 7 0.1 0.2 0 0
queen7_7 49 476 7 7 0.0 7 7 0.0 7 7 0.0 7 7 0.3 7 7 0.2 7 7 0.3 0.5 0 0
queen8_12 96 1368 12 12 0.0 12 12 0.1 12 12 0.1 12 12 4.5 12 12 0.4 12 12 0.4 6.1 0 0
queen8_8 64 728 9 9 3.4 9 9 2.8 9 9 2.6 9 9 1199.1 9 9 132.8 9 9 31.6 4.4 1 24
queen9_9 81 1056 10 10 13.8 10 10 159.2 10 10 921.9 9 11 3600.0 10 10 251.2 10 10 789.0 7.1 21 3600.0
r1000.1 1000 14378 20 20 0.4 20 20 0.4 20 20 0.4 20 20 44.2 20 20 7.9 20 20 5.6 0.8 0 1
@@
wap08a 1870 104176 40 40 461.2 40 40 457.8 40 40 1202.0 39 0 3600.1 40 0 3600.0 40 0 3600.0 3600.0 3600.0 3600.0 555the instance wap08a was solved by the method in seperate experiments using a much higher time limit
will199GPIA 701 6772 7 7 0.1 7 7 0.1 7 7 0.1 7 7 0.5 7 7 0.3 7 7 0.3 2.2 0 1
zeroin.i.1 126 4100 49 49 0.0 49 49 0.0 49 49 0.0 49 49 0.1 49 49 0.1 49 49 0.1 0.1 0 0
zeroin.i.2 157 3541 30 30 0.1 30 30 0.0 30 30 0.1 30 30 0.8 30 30 0.1 30 30 0.1 0.0 0 0
zeroin.i.3 157 3540 30 30 0.1 30 30 0.1 30 30 0.0 30 30 0.8 30 30 0.1 30 30 0.1 0.0 0 0
#solved 98 98 95 72 88 86 73 84 87

### v2

@@ 4.4 Experimental results for the bandwidth coloring problem
set #inst. ASS-S-B POP-S-B POPH-S-B ASS-I-B POP-I-B POPH-I-B DFMM[5]  [^5] The authors did not provide the code, so we used the results as reported in [5]. We remark that the used time limit in the paper was 24 hour compared to 1 hour in our experiments
@@ 5 Conclusion
In this paper, we have revisited the partial-ordering based ILP and SAT formulations for the graph coloring problem and have suggested new models for the bandwidth coloring problem based on partial-ordering models.
Our computational study on the graph coloring problem suggests that all three SAT encodings perform similar, with (POP-S) and (POPH-S) solving slightly more instances (98/143) than (ASS-S) (95/143). This holds true for every timelimit up to 1 hour. Moreover, the SAT encodings solve more instances than the ILP formulations. Compared to state-of-the-art approaches, the tested SAT based approaches solved more instances than the approach based on the set cover ILP formulations and have shown to be particularly advantageous for sparse graphs. Moreover, the tested SAT based encodings also solved more instances than reported in [8] and the same amount of instances as [11]. Specifically, (POP-S) and (POPH-S) have proven to be effective on the wap0- and queen-instances. We also remark that the partial-ordering based encodings are as easy to use as the classical assignment-based encodings. As (POP-S) and (POPH-S) have shown superior performance compared to (ASS-S), an interesting line of research could therefore be to incorporate the encodings into other SAT-based frameworks, such as the method presented in [11].
@@
6 Acknowledgment
We are grateful for the valuable remarks of the referees, which helped to improve the paper.
@@ Appendix A Detailed results of the nine models for the 143 DIMACS instances for the GCP
POP-S POPH-S ASS-S POP-I POPH-I ASS-I EC CLICOL CDCL   [^6] The code provided in the repository produced compile errors on our system, so we used the results of the experiments from [11] which did not contain the R-instances
@@
1-FullIns_3 30 100 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 0.0 0.2 0
1-FullIns_4 93 593 5 5 0.1 5 5 0.1 5 5 0.1 5 5 0.1 5 5 0.2 5 5 0.1 3600.1 0.1 0
1-FullIns_5 282 3247 6 6 0.5 6 6 0.6 6 6 0.5 6 6 361.2 6 6 242.9 6 6 29.4 3600.5 0.3 0
1-Insertions_4 67 232 5 5 1.4 5 5 1.2 5 5 1.0 5 5 71.7 5 5 480.4 5 5 106.9 3600.1 1.9 3600.0
@@
2-FullIns_3 52 201 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.0 0.0 0.5 0
2-FullIns_4 212 1621 6 6 0.0 6 6 0.0 6 6 0.1 6 6 1.8 6 6 0.4 6 6 0.2 3600.1 0.1 0
2-FullIns_5 852 12201 7 7 0.7 7 7 0.6 7 7 0.8 6 7 3600.0 7 7 1899.5 7 7 983.6 3602.5 3.2 20
2-Insertions_3 37 72 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.1 4 4 0.1 4 4 0.1 220.2 0.1 0
@@
3-FullIns_3 80 346 6 6 0.1 6 6 0.1 6 6 0.1 6 6 0.0 6 6 0.0 6 6 0.0 0.0 0.5 0
3-FullIns_4 405 3524 7 7 0.1 7 7 0.1 7 7 0.1 7 7 7.6 7 7 1.4 7 7 0.4 3600.2 0.4 0
3-FullIns_5 2030 33751 8 8 4.1 8 8 5.9 8 8 5.8 7 8 3600.0 7 8 3600.0 7 8 3600.0 3602.9 43.0 0
3-Insertions_3 56 110 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.8 4 4 1.1 4 4 1.4 3600.1 0.2 1
@@
4-FullIns_3 114 541 7 7 0.1 7 7 0.1 7 7 0.0 7 7 0.0 7 7 0.0 7 7 0.0 0.0 0.2 0
4-FullIns_4 690 6650 8 8 0.4 8 8 0.5 8 8 0.6 8 8 60.1 8 8 0.6 8 8 0.5 3600.7 1.2 0
4-FullIns_5 4146 77305 9 9 63.8 9 9 93.3 9 9 75.6 8 9 3600.0 8 9 3600.0 8 9 3600.0 3629.8 678.1 730
4-Insertions_3 79 156 4 4 0.1 4 4 0.1 4 4 0.2 4 4 17.5 4 4 15.7 4 4 48.3 3600.2 0.2 417
@@
anna 138 493 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 0.0 0.4 0
ash331GPIA 662 4181 4 4 0.1 4 4 0.1 4 4 0.1 4 4 6.3 4 4 16.0 4 4 28.6 3611.7 0.2 0
ash608GPIA 1216 7844 4 4 0.1 4 4 0.1 4 4 0.0 4 4 10.7 4 4 34.1 4 4 74.3 3600.0 0.5 4
ash958GPIA 1916 12506 4 4 0.1 4 4 0.1 4 4 0.1 4 4 57.5 4 4 217.9 4 4 872.6 3600.0 1.4 29
david 87 406 11 11 0.1 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.0 0.0 0.0 0
@@
DSJR500.1 500 3555 12 12 0.0 12 12 0.0 12 12 0.0 12 12 1.2 12 12 0.3 12 12 0.2 255.1 0.9 0
DSJR500.1c 500 121275 81 89 3600.0 80 89 3600.0 80 89 3600.0 74 0 3600.1 81 0 3600.0 75 89 3600.3 695.1 52.8 3600.0
DSJR500.5 500 58862 122 131 3600.0 122 131 3600.0 122 131 3600.0 115 0 3600.1 122 122 334.5 122 122 1373.1 2602.2 65.8 3600.0
@@
fpsol2.i.1 269 11654 65 65 0.1 65 65 0.1 65 65 0.1 65 65 1.2 65 65 0.3 65 65 0.3 0.4 0.6 0
fpsol2.i.2 363 8691 30 30 0.1 30 30 0.2 30 30 0.0 28 30 3600.0 30 30 0.2 30 30 0.4 0.4 0.3 0
fpsol2.i.3 363 8688 30 30 0.0 30 30 0.0 30 30 0.0 29 30 3600.0 30 30 0.2 30 30 0.4 0.4 0.1 0
games120 120 638 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 9 9 0.0 0.0 0.1 0
homer 556 1629 13 13 0.0 13 13 0.0 13 13 0.0 13 13 61.2 13 13 0.1 13 13 0.1 0.1 0.30 0
huck 74 301 11 11 0.0 11 11 0.0 11 11 0.0 11 11 0.2 11 11 0.0 11 11 0.0 0.0 0.1 0
inithx.i.1 519 18707 54 54 0.1 54 54 0.0 54 54 0.1 54 54 0.8 54 54 0.3 54 54 0.3 1.3 0.3 0
inithx.i.2 558 13979 31 31 0.1 31 31 0.0 31 31 0.1 29 31 3600.0 31 31 0.2 31 31 0.4 0.2 0.4 0
inithx.i.3 559 13969 31 31 0.1 31 31 0.2 31 31 0.1 29 31 3600.0 31 31 0.2 31 31 0.3 0.3 0.2 0
jean 77 254 10 10 0.0 10 10 0.1 10 10 0.0 10 10 0.0 10 10 0.0 10 10 0.0 0.0 0.1 0
latin_square_10 900 307350 90 132 3600.0 90 132 3600.0 90 132 3600.0 - - - - - - - - - 3638.4 - 3600.0
le450_15a 450 8168 15 15 0.8 15 15 0.3 15 15 0.4 15 16 3600.1 15 15 328.1 15 15 307.6 3633.7 1.1 4
le450_15b 450 8169 15 15 0.7 15 15 0.2 15 15 0.2 15 15 1039.7 15 15 567.0 15 15 293.2 3600.0 0.3 1
le450_15c 450 16680 15 23 3600.0 15 23 3600.0 15 23 3600.0 15 23 3600.0 15 0 3600.0 15 0 3600.7 3600.0 63.6 3600.0
le450_15d 450 16750 15 24 3600.0 15 24 3600.0 15 24 3600.0 15 0 3600.0 15 0 3600.0 15 0 3600.4 3600.0 53.8 3600.0
le450_25a 450 8260 25 25 0.1 25 25 0.1 25 25 0.1 25 25 26.8 25 25 1.3 25 25 0.6 1.8 0.2 0
le450_25b 450 8263 25 25 0.0 25 25 0.0 25 25 0.0 25 25 10.9 25 25 0.8 25 25 0.6 1.9 0.2 0
Table 4: Results of the nine models for the 134 DIMACS instances and the 9 R-instances for the GCP
POP-S POPH-S ASS-S POP-I POPH-I ASS-I EC CLICOL CDCL
@@
le450_5a 450 5714 5 5 0.1 5 5 0.0 5 5 0.0 5 5 108.5 5 5 23.2 5 5 62.9 3600.0 0.1 41
le450_5b 450 5734 5 5 0.0 5 5 0.0 5 5 0.0 5 5 328.8 5 5 41.2 5 5 47.6 3600.0 0.1 9
le450_5c 450 9803 5 5 0.0 5 5 0.1 5 5 0.0 5 5 92.4 5 5 47.5 5 5 19.3 3600.0 0.1 2
le450_5d 450 9757 5 5 0.1 5 5 0.1 5 5 0.0 5 5 19.8 5 5 49.8 5 5 46.8 3584.7 0.1 2
miles1000 128 3216 42 42 0.1 42 42 0.1 42 42 0.1 42 42 0.5 42 42 0.1 42 42 0.1 0.1 0.6 0
miles1500 128 5198 73 73 0.1 73 73 0.2 73 73 0.2 73 73 0.5 73 73 0.3 73 73 0.3 0.1 0.1 0
miles250 125 387 8 8 0.1 8 8 0.0 8 8 0.0 8 8 0.0 8 8 0.0 8 8 0.0 0.0 0.1 0
miles500 128 1170 20 20 0.1 20 20 0.0 20 20 0.1 20 20 0.0 20 20 0.0 20 20 0.0 0.0 0.1 0
miles750 128 2113 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.1 31 31 0.1 31 31 0.1 0.0 0.1 0
mug100_1 100 166 4 4 0.1 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.1 4 4 0.2 0.6 0.1 0
mug100_25 100 166 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.4 4 4 0.2 0.5 0.1 0
mug88_1 88 146 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.1 4 4 0.1 4 4 0.2 0.3 0.1 0
mug88_25 88 146 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.2 4 4 0.1 4 4 0.2 0.3 0.1 0
mulsol.i.1 138 3925 49 49 0.0 49 49 0.1 49 49 0.1 49 49 0.4 49 49 0.1 49 49 0.1 0.1 0.1 0
mulsol.i.2 173 3885 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.3 31 31 0.1 31 31 0.1 0.0 0.1 0
mulsol.i.3 174 3916 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.2 31 31 0.1 31 31 0.1 0.0 0.2 0
mulsol.i.4 175 3946 31 31 0.0 31 31 0.0 31 31 0.0 31 31 0.4 31 31 0.1 31 31 0.1 0.0 0.1 0
mulsol.i.5 176 3973 31 31 0.0 31 31 0.0 31 31 0.0 30 31 3600.0 31 31 0.1 31 31 0.1 0.0 0.1 0
myciel3 11 20 4 4 0.0 4 4 0.1 4 4 0.0 4 4 0.0 4 4 0.0 4 4 0.0 0.0 0.1 0
myciel4 23 71 5 5 0.0 5 5 0.0 5 5 0.0 5 5 0.1 5 5 0.1 5 5 0.1 4.8 0.1 0
myciel5 47 236 6 6 0.2 6 6 0.2 6 6 0.2 6 6 37.5 6 6 42.0 6 6 44.3 3600.0 0.5 0
myciel6 95 755 7 7 99.4 7 7 74.3 7 7 63.6 6 7 3600.0 6 7 3600.0 5 7 3600.0 3600.1 1045.3 0
myciel7 191 2360 6 8 3600.0 6 8 3600.0 6 8 3600.0 5 8 3600.0 5 8 3600.0 5 8 3600.0 3600.7 3600.0 0
@@
qg.order30 900 26100 30 30 0.8 30 30 0.4 30 30 0.5 30 35 3600.2 30 30 77.6 30 30 186.5 3600.0 4.5 0
qg.order40 1600 62400 40 40 7.5 40 40 1.9 40 40 2.4 40 0 3600.0 40 40 1798.7 40 43 3600.2 3600.0 256.4 8
qg.order60 3600 212400 60 60 1813.9 60 60 835.9 60 60 28.8 60 0 3600.1 60 62 3601.5 60 0 3600.0 3600.0 315.9 347
@@
queen5_5 25 160 5 5 0.0 5 5 0.0 5 5 0.1 5 5 0.0 5 5 0.0 5 5 0.0 0.0 0.4 0
queen6_6 36 290 7 7 0.1 7 7 0.1 7 7 0.0 7 7 0.4 7 7 0.2 7 7 0.1 0.2 0.1 0
queen7_7 49 476 7 7 0.0 7 7 0.0 7 7 0.0 7 7 0.3 7 7 0.2 7 7 0.3 0.5 0.1 0
queen8_12 96 1368 12 12 0.0 12 12 0.1 12 12 0.1 12 12 4.5 12 12 0.4 12 12 0.4 6.1 0.3 0
queen8_8 64 728 9 9 3.4 9 9 2.8 9 9 2.6 9 9 1199.1 9 9 132.8 9 9 31.6 4.4 13.9 1
queen9_9 81 1056 10 10 13.8 10 10 159.2 10 10 921.9 9 11 3600.0 10 10 251.2 10 10 789.0 7.1 3600.0 21
r1000.1 1000 14378 20 20 0.4 20 20 0.4 20 20 0.4 20 20 44.2 20 20 7.9 20 20 5.6 0.8 0.6 0
@@
wap08a 1870 104176 40 40 461.2 40 40 457.8 40 40 1202.0 39 0 3600.1 40 0 3600.0 40 0 3600.0 3600.0 3600.0 3600.0
will199GPIA 701 6772 7 7 0.1 7 7 0.1 7 7 0.1 7 7 0.5 7 7 0.3 7 7 0.3 2.2 0.8 0
zeroin.i.1 126 4100 49 49 0.0 49 49 0.0 49 49 0.0 49 49 0.1 49 49 0.1 49 49 0.1 0.1 0.1 0
zeroin.i.2 157 3541 30 30 0.1 30 30 0.0 30 30 0.1 30 30 0.8 30 30 0.1 30 30 0.1 0.0 0.1 0
zeroin.i.3 157 3540 30 30 0.1 30 30 0.1 30 30 0.0 30 30 0.8 30 30 0.1 30 30 0.1 0.0 0.1 0
#solved 98 98 95 72 88 86 73 98 84